OpenStudy (anonymous):

I have $28 n bank acct and save 18.25 a week. my friend has $161 and withdraws 15.00 a week. when will balances be the same. I can get the answer but dont know the correct equation for it.

OpenStudy (anonymous):

28 + 18.25x = y 161 - 15.00x+ y

OpenStudy (anonymous):

where x is the number of days and y is the final value in your account. In this case we want the y to be equal so we can join both equations

OpenStudy (anonymous):

28 + 18.25x = 161 - 15.00x Sorry on my first post I made a mistake its 161 - 15.00x= y

OpenStudy (anonymous):

you have one equation and 1 unknown. If you can't solve from my last equation just ask for more help.

OpenStudy (anonymous):

That doesn't sound right.

OpenStudy (anonymous):

thx Romero now maybe I can get the next two on my own

OpenStudy (anonymous):

28 + 18.25w = 161 - 15w (18.25 +15)w = 161 - 28 -> w = 4 wks

OpenStudy (anonymous):

I didn't want to do the simplification because it's easy to make mistakes like cholrophyll did.

OpenStudy (anonymous):

Yep, I'm alway have to fix my post right when I see it clear :")